OpenStudy (anonymous):

Solve for x: |x + 9| = 17

OpenStudy (theviper):

x+9=17 x+9=-17 solve for both :)

Parth (parthkohli):

The first step is fetching two different equations:\[x + 9 = +17\\x + 9 = -17 \]

Parth (parthkohli):

Individually solve, and done.

OpenStudy (theviper):

Got it @heartbreak ??